OpenStudy (anonymous):

Twice a number decreased by two is equivalent to than number increased by five

OpenStudy (anonymous):

2x-2=x+5

OpenStudy (anonymous):

SHOULD I SOLVE EVEN IT DIDNT SAY SO

OpenStudy (anonymous):

that is the solution, your translating terms.

OpenStudy (anonymous):

K GREAT THNX CHARLIE

OpenStudy (anonymous):

2x-2=x+5 2x-x=5+2 x=7